For the given question below:Suppose an attractive nuclear force acts between two protons which may be written as F= C e^(-ke) /r^2. a) Write down the dimensional formulae and appropriate SI units of C and k.b) Suppose that k= 1fermi^-1 and that the repulsive electric force between the protons is just balanced by the attractive nuclear force when the separation is 5 fermi. Find the value of C.Please someone answer part (b).And For part a) why is e^(-kr) a pure number.

Grade:

1 Answers

Arun
5 years ago
Expression of electrical force F = C * e^-kr/r^2 Since e^-kr is a pure number. So, dimecsional formulae of F = dim ensional formulae of C/dim ensional formulae of r^2 Or, [MLT^–2][L^2] = dimensional formulae of C = [ML^3T^–2] Unit of C = unit of force × unit of r^2 = Newton × m^2 = Newton–m^2 Since –kr is a number hence dimensional formulae of K = 1/dim entional formulae of r = [L^-1] Unit of k = m^-1